OpenStudy (anonymous):

Sove by factoring. 11x^3 -7x^2=0

OpenStudy (mathstudent55):

Do you see a common factor for the two terms?

OpenStudy (anonymous):

x squared??

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Right. Now factor x^2 from both terms.

OpenStudy (anonymous):

Oh wow lol that was a lot simpler than I thought, was over thinking it I guess

OpenStudy (mathstudent55):

Can you finish now?

OpenStudy (anonymous):

x=0 and x=7/11.

OpenStudy (anonymous):

Okay I understand how you got 7/11 but I don't understand how you got 0...

OpenStudy (mathstudent55):

The factoring is: x^2(11x - 7) = 0 A product equals zero if one or more factors equal zero. So you set each factor equal to zero to find x. x^2 = 0 or 11x - 7 = 0 x = 0 or 11x = 7 x = 0 or x = 7/11

OpenStudy (anonymous):

Ohh.... I get it now Thanks! I haven't done factoring since September and I need to do well on my final lol.

OpenStudy (mathstudent55):

You're welcome. Good luck.

OpenStudy (anonymous):

Thanks